# Simultaneous Khintchine theorem on manifolds in positive characteristics: convergence case

### v1

@@ 3. Preliminaries in Geometry of Numbers
Recall that for $\ell\in\mathbb{N}$, $\mathcal{K}_{\ell}=\mathcal{K}(T^{1/\ell})$ and $\mathcal{R}=\mathbb{F}_{q}[T]$. Note that $\mathcal{K}_{\ell}=\mathcal{K}$ when $\ell=1$. For $i=1,\dots,n+1$ and $g\in\operatorname{SL}(n+1,\mathcal{K}_{\ell})$, we define the $i$-th successive minimum of the discrete subgroup $g\mathcal{R}^{n+1}\subseteq\mathcal{K}_{\ell}^{n+1}$ as
$\lambda_{i}(g\mathcal{R}^{n+1}):=\inf\left\{\lambda>0:B(0,\lambda)\cap g\mathcal{R}^{n+1}\text{ contains }i\text{ linearly independent vectors over }\mathcal{K}_{\ell}\right\},$
@@
We define $G:=\operatorname{SL}(n+1,\mathcal{K})$ and $\Gamma:=\operatorname{SL}(n+1,\mathcal{R})$. Then the associated homogeneous space $\mathcal{X}_{n+1}:=G/\Gamma$ can be identified with the space of all unimodular lattices in $\mathcal{K}^{n+1}$. Now we recall the function field analogue of Minkowski’s second theorem from [Mah41, Equations (24) and (25)].
@@
$\Lambda^{*}:=\left\{\mathbf{y}\in\mathcal{K}^{n+1}:\forall\mathbf{x}\in\Lambda,\langle\mathbf{x},\mathbf{y}\rangle\in\mathcal{R}\right\},$
where $\langle\mathbf{x},\mathbf{y}\rangle=\sum_{i=1}^{n}x_{i}y_{i}$. We first state the duality theorem on $\mathcal{K}^{n+1}$, originally due to Mahler[Mah41, Equation (28)] (see also [BK25, Lemma 5.4]).
@@
and $g^{*}$ is defined as $g^{*}:=(g^{t})^{-1}$. We prove the following generalization of Theorem 3.2 for some specific discrete subgroups in $\mathcal{K}_{\ell}^{n+1}$.
@@
Let $g\in\operatorname{GL}(n+1,\mathcal{K}_{\ell})$ be of the form (3.3) and $\Lambda$ be a lattice in $\mathcal{K}^{n+1}$. Then
$q^{-\frac{\ell-1}{\ell}}\leq\lambda_{i}(g^{*}\Lambda^{*})\lambda_{n+2-i}(g\Lambda)\leq q^{\frac{\ell-1}{\ell}},~\forall i=1,\dots,n+1.$ (3.4)
@@
Without any loss of generality, we may assume that $a_{i}\in\left\{0,1,\dots,\ell-1\right\}$; for, if necessary, we may replace $\Lambda$ by the lattice $\operatorname{diag}\left(T^{\left\lfloor\frac{a_{1}}{\ell}\right\rfloor},\dots,T^{\left\lfloor{\frac{a_{n+1}}{\ell}}\right\rfloor}\right)\Lambda$, where $\lfloor c\rfloor:=\max\{k\in\mathbb{Z}:k\leq c\}$ for $c\in\mathbb{R}.$ Note that,
@@
$q^{-\frac{\ell-1}{\ell}}\|\mathbf{v}\|\leq\|g\|^{-1}\|\mathbf{v}\|=\|g\|^{-1}\|gg^{-1}\mathbf{v}\|\leq\|g^{-1}\mathbf{v}\|\leq\|\mathbf{v}\|,\forall\mathbf{v}\in\mathcal{K}^{n+1}.$ (3.5)
Since $\mathcal{K}_{\ell}$ is a finite extension over $\mathcal{K}$, then $\mathbf{v}_{1},\dots,\mathbf{v}_{n+1}\in\Lambda$ are linearly independent over $\mathcal{K}$ if and only if $\mathbf{v}_{1},\dots,\mathbf{v}_{n+1}$ are linearly independent over $\mathcal{K}_{\ell}$. Hence,
@@
$\lambda_{i}(g\Lambda)\leq q^{\frac{\ell-1}{\ell}}\lambda_{i}(\Lambda),~\forall i=1,\dots,n+1.$ (3.7)
@@
$\lambda_{i}(g^{-1}\Lambda)\leq\lambda_{i}(\Lambda),~\forall i=1,\dots,n+1.$ (3.10)
@@
$q^{-\frac{\ell-1}{\ell}}\lambda_{i}(\Lambda^{\ast})\leq\lambda_{i}(g^{\ast}\Lambda^{\ast})\leq\lambda_{i}(\Lambda^{\ast}),~\forall i=1,\dots,n+1,$ (3.11)
since $g$ is diagonal, so that $g^{-1}$ is the same as its transpose. Multiplying inequalities in (3.7) with j=n+2ij=n+2-i and (3.11) now yields that
$\lambda_{i}(g^{\ast}\Lambda^{\ast})\lambda_{n+2-i}(g\Lambda)\leq q^{\frac{\ell-1}{\ell}}\lambda_{i}(\Lambda^{\ast})\lambda_{n+2-i}(\Lambda),~\forall i=1,\dots,n+1.$
@@
$q^{-\frac{\ell-1}{\ell}}\lambda_{i}(\Lambda^{\ast})\lambda_{n+2-i}(\Lambda)\leq\lambda_{i}(g^{\ast}\Lambda^{\ast})\lambda_{n+2-i}(g\Lambda)\leq q^{\frac{\ell-1}{\ell}}\lambda_{i}(\Lambda^{\ast})\lambda_{n+2-i}(\Lambda),~\forall i=1,\dots,n+1.$
@@
Recall from [Mah41] that convex bodies in $\mathcal{K}^{n+1}$ are precisely the sets of the form $g\mathcal{O}_{\mathcal{K}}^{n+1}$, where $g\in\operatorname{GL}(n+1,\mathcal{K}_{\ell})$. Then, clearly $\operatorname{Vol}(g\mathcal{O}_{\mathcal{K}}^{n+1})=|\det(g)|$. We now state the following lemma, due to Bagshaw and Kerr [BK25], that counts the number of lattice points in a convex body:
@@
We generalize the above result for specific discrete subgroups of the form $g\Lambda\subseteq\mathcal{K}_{\ell}^{n+1}$ with $g$ being diagonal over $\mathcal{K}_{\ell}$ and $\Lambda$ being a lattice in $\mathcal{K}$ under the assumption $\lambda_{n+1}(g\Lambda)\leq q^{c}$ for some $c>0$.
@@
Let $g=\operatorname{diag}\left(T^{a_{1}+\frac{\alpha_{1}}{\ell}},\dots,T^{a_{n+1}+\frac{\alpha_{n+1}}{\ell}}\right)\in\operatorname{GL}(n+1,\mathcal{K}_{\ell})$, where $a_{i}\in\mathbb{Z}$ and $\alpha_{i}\in\{0,1,\dots,(\ell-1)\}$. Consider any $c\geq 0$ such that $\lambda_{n+1}(g\Lambda)\leq q^{c}$, where $\Lambda=u\mathcal{R}^{n+1}$ and $u\in\operatorname{GL}(n+1,\mathcal{K})$. Then,
@@
Let $g\mathbf{v}_{1},\dots,g\mathbf{v}_{n+1}$ be linearly independent vectors in $g\Lambda$ with $\|g\mathbf{v}_{i}\|=\lambda_{i}(g\Lambda)$. Clearly, $g]\mathbf{v}_{1},\dots,[g]\mathbf{v}_{n+1}$ are linearly independent. Then it follows from (3.12) and the hypothesis $\lambda_{n+1}(g\Lambda)\leq q^{c}$ that $\lambda_{n+1}([g]\Lambda)\leq q^{c}$. Furthermore, (3.12) yields that the parallelepiped spanned by the vectors $g]\mathbf{v}_{1},\dots,[g]\mathbf{v}_{n+1}$ inside $\mathcal{K}^{n+1}$ is actually contained in $T^{\lfloor{c}\rfloor}\mathcal{O}_{\mathcal{K}}^{n+1}$, where $\lfloor c\rfloor=\max\{n\in\mathbb{Z}:n\leq c\}$. Now, in view of (3.12), applying Lemma 3.5, we arrive at the following:
$\begin{split}\#\left\{\mathbf{v}\in\{g\}[g]\Lambda:\|\mathbf{v}\|\leq q^{c}\right\}\leq\#\{\mathbf{v}\in[g]\Lambda:\|\mathbf{v}\|\leq q^{\lfloor{c}\rfloor}\}\leq\frac{q^{(n+1)(c+1)}}{|\det([g])|\cdot|\det(u)|}\leq\frac{q^{(n+1)(c+1)}}{|\det(g)|\cdot|\det(u)|}.\end{split}$
@@ 4.1. Some Preliminary Estimates
In this subsection, we provide several estimates that will be used to prove Theorems 1.2 and 1.3. Recall that the manifold $\mathscr{M}$ is defined by the map $\mathbf{f}:U\subseteq\mathcal{K}^{d}\to\mathcal{K}^{n}$, where
$\mathbf{f}:=(x_{1},\ldots,x_{d},f_{1}(\mathbf{x}),\ldots,f_{m}(\mathbf{x}))=(\mathbf{x},\bm{f}(\mathbf{x})),\,\,\text{for}\,\,\mathbf{x}=(x_{1},\dots,x_{d})\in U$
with $d=\dim(\mathscr{M})$, $m=\operatorname{\text{codim}}(\mathscr{M})$, and $U\subseteq\mathcal{K}^{d}$ being an open set. We also assume that the defining map $\mathbf{f}$ is analytic.
@@
  2. (2)
    $|Qf_{j}(\mathbf{x})-\sum_{i=1}^{d}\partial_{i}f_{j}(\mathbf{x})(Qx_{i}-P_{i}+\Theta_{i})-P_{d+j}+\Theta_{d+j}|<\max\{1,M\}q^{-s}$ for every $j=1,\dots,m$.
@@
(1) is easily seen by considering the $i$-th coordinate of $\mathbf{f}(\mathbf{x})-\mathbf{P}-\bm{\Theta}$ for $i=1,\dots,d$, and using $|Q|=q^{t}$. To prove (2), using the ultrametric inequality, we obtain
$\bigg|Qf_{j}(\mathbf{x})-\sum_{i=1}^{d}\partial_{i}f_{j}(\mathbf{x})\big(Qx_{i}-P_{i}-\Theta_{i}\big)-P_{d+j}-\Theta_{d+j}\bigg|$ $\leq\max\Big\{|Qf_{j}(\mathbf{x})-P_{d+j}-\Theta_{d+j}|,$
$\max_{i=1,\dots,d}|\partial_{i}f_{j}(\mathbf{x})|\cdot|Qx_{i}-P_{i}-\Theta_{i}|\Big\}.$
Now, from the conditions $\big\|\mathbf{f}(\mathbf{x})-\tfrac{\mathbf{P}+\bm{\Theta}}{Q}\big\|<q^{-(s+t)}\,\,\text{and}\,\,|Q|=q^{t}$, it follows that
$|Qf_{j}(\mathbf{x})-P_{d+j}-\Theta_{d+j}|<q^{-s}.$
Combining this and (1), we conclude
$\max\Big\{|Qf_{j}(\mathbf{x})-P_{d+j}-\Theta_{d+j}|,\max_{i=1,\dots,d}|\partial_{i}f_{j}(\mathbf{x})|\cdot|Qx_{i}-P_{i}-\Theta_{i}|\Big\}<\max\{1,M\}q^{-s}.$
@@
Suppose that $\mathbf{x}\in U$ and $(\mathbf{P},Q)\in\mathcal{R}^{n}\times(\mathcal{R}\setminus\{0\})$ satisfy the hypothesis of Lemma 4.1. Let $\mathbf{x}\in B(\mathbf{x}_{0},q^{-\frac{s+t}{2}})$ for some $\mathbf{x}_{0}=(x_{1,0},\cdots,x_{d,0})\in U$, then,
  1. (1)
    $|Qx_{i,0}-P_{i}-\Theta_{i}|<\max\{q^{-s},q^{t-\frac{s+t}{2}}\}$ for $i=1,\dots,d$.
  2. (2)
@@
    $\left|Qf_{j}(\mathbf{x}_{0})-\sum_{i=1}^{d}\partial_{i}f_{j}(\mathbf{x}_{0})(Qx_{0,i}-P_{i}-\Theta_{i})-P_{d+j}-\Theta_{d+j}\right|\leq\max\{1,M\}q^{-s}.$
@@
Let $\mathbf{x}_{0}=(x_{1,0},\cdots,x_{d,0})\in U$. By Lemma 4.1, we have the vector $\mathbf{x}$ satisfies (1) and (2). Hence,
$|Qx_{i,0}-P_{i}-\Theta_{i}|$ $\leq$ $\max\left\{|Qx_{i}-P_{i}-\Theta_{i}|,|Q|\cdot|x_{i}-x_{i,0}|\right\}$
$\leq$ $\max\{q^{-s},q^{t-\frac{s+t}{2}}\}.$
Note that by Lemma 2.2, we have
$\begin{split}\left|Qf_{j}(\mathbf{x}_{0})-\sum_{i=1}^{d}\partial_{i}f_{j}(\mathbf{x}_{0})(Qx_{0,i}-P_{i}+\Theta_{i})-P_{d+j}-\Theta_{d+j}\right|\\ \leq\max\Bigg\{\left|Qf_{j}(\mathbf{x})-\sum_{i=1}^{d}\partial_{i}f_{j}(\mathbf{x})(Qx_{i}-P_{i}-\Theta_{i})-P_{d+j}-\Theta_{d+j}\right|,\\ |Q|\cdot\left|f_{j}(\mathbf{x}_{0})-f_{j}(\mathbf{x})-\sum_{i=1}^{d}\partial_{i}f_{j}(\mathbf{x})(x_{0,i}-x_{i})\right|,\\ \max_{i=1,\dots,d}|\partial_{i}f_{j}(\mathbf{x}_{0})-\partial_{i}f_{j}(\mathbf{x})|\cdot|Qx_{i}-P_{i}-\Theta_{i}|\Bigg\}\\ \leq\max\{1,M\}q^{-s}.\end{split}$
4.2. Analysis on Generic Part
We first define the generic part. To do so, we first define the diagonal flow $g_{s,t}$ as
$g_{s,t}:=\operatorname{diag}\left\{\underbrace{T^{\frac{(d+2)t+2s}{2(n+1)}},\dots,T^{\frac{(d+2)t+2s}{2(n+1)}}}_{m},\overbrace{T^{\frac{(d+2)t+2s}{2(n+1)}-\frac{s+t}{2}},\dots,T^{\frac{(d+2)t+2s}{2(n+1)}-\frac{s+t}{2}}}^{d},T^{\frac{(d+2)t+2s}{2(n+1)}-s-t}\right\}.$
Note that the entries of the diagonal flow $g_{s,t}$ come from the extended field $\mathcal{K}_{2(n+1)}$. For $\mathbf{x}\in U$, denote by $J(\mathbf{x})$ the Jacobian matrix of $\mathbf{f}$ at $\mathbf{x}$, i.e.,
$J(\mathbf{x}):=\left(\frac{\partial f_{i}}{\partial x_{j}}(\mathbf{x})\right)_{1\leq i\leq d,1\leq j\leq m}.$

### v2

@@ 3. Preliminaries in Geometry of Numbers
Recall that for $\ell\in\mathbb{N}$, $\mathcal{K}_{\ell}=\mathcal{K}(T^{1/\ell})$ and $\mathcal{R}=\mathbb{F}_{q}[T]$. Note that $\mathcal{K}_{\ell}=\mathcal{K}$ when $\ell=1$. For $i=1,\dots,n+1$ and $g\in\operatorname{SL}(n+1,\mathcal{K}_{\ell})$, we define the $i$-th successive minimum of the discrete subgroup $g\mathcal{R}^{n+1}\subseteqq\mathcal{K}_{\ell}^{n+1}$ as
$\lambda_{i}(g\mathcal{R}^{n+1}){\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\inf\left\{\lambda>0:B(0,\lambda)\cap g\mathcal{R}^{n+1}\text{ contains }i\text{ linearly independent vectors over }\mathcal{K}_{\ell}\right\},$
@@
We define $G{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\operatorname{SL}(n+1,\mathcal{K})$ and $\Gamma{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\operatorname{SL}(n+1,\mathcal{R})$. Then the associated homogeneous space $\mathcal{X}_{n+1}{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}G/\Gamma$ can be identified with the space of all unimodular lattices in $\mathcal{K}^{n+1}$. Now we recall the function field analog of Minkowski’s second theorem from [Mah41, Equations (24) and (25)].
@@
$\Lambda^{*}{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\left\{\mathbf{y}\in\mathcal{K}^{n+1}:\forall\mathbf{x}\in\Lambda,\langle\mathbf{x},\mathbf{y}\rangle\in\mathcal{R}\right\},$
where $\langle\mathbf{x},\mathbf{y}\rangle=\sum_{i=1}^{n+1}x_{i}y_{i}$. We first state the duality theorem on $\mathcal{K}^{n+1}$, originally due to Mahler [Mah41, Equation (28)] (see also [BK25, Lemma 5.4]).
@@
and $g^{*}$ is defined as $g^{*}{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}(g^{t})^{-1}$. We prove the following generalization of Theorem 3.2 for some specific discrete subgroups in $\mathcal{K}_{\ell}^{n+1}$.
@@
Let $g\in\operatorname{GL}(n+1,\mathcal{K}_{\ell})$ be of the form (3.3), and let $\Lambda$ be a lattice in $\mathcal{K}^{n+1}$. Then
$q^{-\frac{\ell-1}{\ell}}\leq\lambda_{i}(g^{*}\Lambda^{*})\lambda_{n+2-i}(g\Lambda)\leq q^{\frac{\ell-1}{\ell}},\penalty 10000\ \forall i=1,\dots,n+1.$ (3.4)
@@
Without any loss of generality, we may assume that $a_{i}\in\left\{0,1,\dots,\ell-1\right\}$; for if necessary, we may replace $\Lambda$ with the lattice $\operatorname{diag}\left(T^{\left\lfloor\frac{a_{1}}{\ell}\right\rfloor},\dots,T^{\left\lfloor{\frac{a_{n+1}}{\ell}}\right\rfloor}\right)\Lambda$, where $\lfloor c\rfloor{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\max\{k\in\mathbb{Z}:k\leq c\}$ for $c\in\mathbb{R}$.
Since $a_{i}\in\{0,1,\dots,\ell-1\}$, it follows that
@@
$q^{-\frac{\ell-1}{\ell}}\|\mathbf{v}\|\leq\|g\|^{-1}\|\mathbf{v}\|=\|g\|^{-1}\|gg^{-1}\mathbf{v}\|\leq\|g^{-1}\mathbf{v}\|\leq\|\mathbf{v}\|,\forall\mathbf{v}\in\mathcal{K}^{n+1},$ (3.5)
where, for any diagonal matrix ${\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\operatorname{diag}(\beta_{1},\dots,\beta_{n+1})\in\operatorname{Mat}_{n+1}(\mathcal{K}_{\ell}), D\|D\|$ is defined by $\max_{1\leq i\leq n+1}|a_{i}|$. Since $\mathcal{K}_{\ell}$ is a finite extension over $\mathcal{K}$, $\mathbf{v}_{1},\dots,\mathbf{v}_{n+1}\in\Lambda$ are linearly independent over $\mathcal{K}$ if and only if $\mathbf{v}_{1},\dots,\mathbf{v}_{n+1}$ are linearly independent over $\mathcal{K}_{\ell}$. Hence,
@@
$\lambda_{i}(g\Lambda)\leq q^{\frac{\ell-1}{\ell}}\lambda_{i}(\Lambda),\penalty 10000\ \forall i=1,\dots,n+1.$ (3.7)
@@
$\lambda_{i}(g^{-1}\Lambda)\leq\lambda_{i}(\Lambda),\penalty 10000\ \forall i=1,\dots,n+1.$ (3.10)
@@
$q^{-\frac{\ell-1}{\ell}}\lambda_{i}(\Lambda^{\ast})\leq\lambda_{i}(g^{\ast}\Lambda^{\ast})\leq\lambda_{i}(\Lambda^{\ast}),\penalty 10000\ \forall i=1,\dots,n+1,$ (3.11)
since $g$ is diagonal, so that $g^{-1}$ is the same as its transpose. Multiplying inequalities in (3.7) with $i$ replaced by n+2in+2-i and (3.11) now yields that
$\lambda_{i}(g^{\ast}\Lambda^{\ast})\lambda_{n+2-i}(g\Lambda)\leq q^{\frac{\ell-1}{\ell}}\lambda_{i}(\Lambda^{\ast})\lambda_{n+2-i}(\Lambda),\penalty 10000\ \forall i=1,\dots,n+1.$
@@
$q^{-\frac{\ell-1}{\ell}}\lambda_{i}(\Lambda^{\ast})\lambda_{n+2-i}(\Lambda)\leq\lambda_{i}(g^{\ast}\Lambda^{\ast})\lambda_{n+2-i}(g\Lambda)\leq q^{\frac{\ell-1}{\ell}}\lambda_{i}(\Lambda^{\ast})\lambda_{n+2-i}(\Lambda),\penalty 10000\ \forall i=1,\dots,n+1.$
@@
Recall from [Mah41] that convex bodies in $\mathcal{K}^{n+1}$ are precisely the sets of the form $g\mathcal{O}_{\mathcal{K}}^{n+1}$, where $g\in\operatorname{GL}(n+1,\mathcal{K}_{\ell})$. Then, clearly $\operatorname{Vol}(g\mathcal{O}_{\mathcal{K}}^{n+1})=|\det(g)|$. We now state the following lemma, due to Bagshaw and Kerr [BK25], which counts the number of lattice points in a convex body:
@@
We generalize the above result for specific discrete subgroups of the form $g\Lambda\subseteqq\mathcal{K}_{\ell}^{n+1}$ with $g$ being diagonal over $\mathcal{K}_{\ell}$ and $\Lambda$ being a lattice in $\mathcal{K}$ under the assumption $\lambda_{n+1}(g\Lambda)\leq q^{c}$ for some $c>0$. To proceed, we first need to establish the following straightforward lemma:
@@
Let $n\in\mathbb{N}$ and $\Lambda\subseteqq\mathcal{K}^{n+1}$ be a lattice. Suppose that $\mathbf{v}_{1},\dots,\mathbf{v}_{n+1}$ are linearly independent vectors in $\Lambda$. Then there exists a basis $\mathbf{w}_{1},\dots,\mathbf{w}_{n+1}$ of $\Lambda$ such that
$\mathbf{v}_{1}=a_{11}\mathbf{w}_{1},\mathbf{v}_{2}=a_{21}\mathbf{w}_{1}+a_{22}\mathbf{w}_{2},\dots,\mathbf{v}_{n+1}=a_{(n+1)1}\mathbf{w}_{1}+\dots+a_{(n+1)(n+1)}\mathbf{w}_{n+1},$
where $a_{ij}$s are from $\mathcal{R}$, satisfying the following:
$a_{ii}\neq 0,|a_{ii}|\geq|a_{ij}|,\text{ for }i=1,\dots,n+1,j=1,\dots,i.$ (3.12)
Lemma 3.6 can easily be proved by adapting the argument given in the proofs of [Cas97, Theorem 1 and Corollary 1, §1.2]. As the argument therein is quite elementary, we leave the details to the reader. However, the conditions mentioned above in (3.12) are important for us. A simple induction argument shows that
$\{\mathbf{w}_{1},\dots,\mathbf{w}_{n+1}\}\subseteqq\mathcal{O}_{\mathcal{K}}\mathbf{v}_{1}+\mathcal{O}_{\mathcal{K}}\mathbf{v}_{2}+\dots+\mathcal{O}_{\mathcal{K}}\mathbf{v}_{n+1}.$
Consequently, the parallelepiped spanned by the vectors $\mathbf{v}_{1},\dots,\mathbf{v}_{n+1}$ contains a fundamental domain of the lattice $\Lambda$.
Lemma 3.7.
Let $g=\operatorname{diag}\left(T^{a_{1}+\frac{\alpha_{1}}{\ell}},\dots,T^{a_{n+1}+\frac{\alpha_{n+1}}{\ell}}\right)\in\operatorname{GL}(n+1,\mathcal{K}_{\ell})$, where $a_{i}\in\mathbb{Z}$ and $\alpha_{i}=0,1,\dots,\ell-1$. Consider any $c\in\mathbb{R}$ such that $\lambda_{n+1}(g\Lambda)\leq q^{c}$, where $\Lambda=u\mathcal{R}^{n+1}$ and $u\in\operatorname{GL}(n+1,\mathcal{K})$. Then,
@@
Let $g\mathbf{v}_{1},\dots,g\mathbf{v}_{n+1}$ be linearly independent vectors in $g\Lambda$ with $\|g\mathbf{v}_{i}\|=\lambda_{i}(g\Lambda)$. Clearly, $g]\mathbf{v}_{1},\dots,[g]\mathbf{v}_{n+1}$ are linearly independent. Then it follows from (3.13) and the hypothesis $\lambda_{n+1}(g\Lambda)\leq q^{c}$ that $\lambda_{n+1}([g]\Lambda)\leq q^{c}$. Furthermore, (3.13) yields that the parallelepiped spanned by the vectors $g]\mathbf{v}_{1},\dots,[g]\mathbf{v}_{n+1}$ inside $\mathcal{K}^{n+1}$ is actually contained in $T^{\lfloor{c}\rfloor}\mathcal{O}_{\mathcal{K}}^{n+1}$, where $\lfloor c\rfloor=\max\{n\in\mathbb{Z}:n\leq c\}$. As a consequence, $T^{\lfloor{c}\rfloor}\mathcal{O}_{\mathcal{K}}^{n+1}$ contains a fundamental domain of $\Lambda$. Thus, by (3.13) and Lemma 3.5,
$\left\{\mathbf{v}\in g\Lambda:\|\mathbf{v}\|\leq q^{c}\right\}\leq\#\{\mathbf{v}\in[g]\Lambda:\|\mathbf{v}\|\leq q^{\lfloor{c}\rfloor}\}\leq\frac{q^{(n+1)(c+1)}}{|\det([g])|\cdot|\det(u)|}\leq\frac{q^{(n+1)(c+1)}}{|\det(g)|\cdot|\det(u)|}.$
@@ 4.1. Some Preliminary Estimates
In this subsection, we provide several estimates that will be used to prove Theorems 1.2 and 1.4. Recall that the manifold $\mathscr{M}$ is defined by the map $\mathbf{f}:U\subseteqq\mathcal{K}^{d}\to\mathcal{K}^{n}$, where
$\mathbf{f}{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}(x_{1},\ldots,x_{d},f_{1}(\mathbf{x}),\ldots,f_{m}(\mathbf{x}))=(\mathbf{x},\bm{f}(\mathbf{x})),\,\,\text{for}\,\,\mathbf{x}=(x_{1},\dots,x_{d})\in U$
with $d=\dim\mathscr{M}$, $m=\operatorname{\text{codim}}\mathscr{M}$, and $U\subseteqq\mathcal{K}^{d}$ being an open set. We also assume that the defining map $\mathbf{f}$ is analytic.
@@
  2. ($\mathfrak{E}$.2)
    $|Qf_{j}(\mathbf{x})-\sum_{i=1}^{d}\partial_{i}f_{j}(\mathbf{x})(Qx_{i}-P_{i}-\Theta_{i})-P_{d+j}-\Theta_{d+j}|<Mq^{-s}$ for every $j=1,\dots,m$.
@@
($\mathfrak{E}$.1) is easily seen by considering the $i$-th coordinate of $\mathbf{f}(\mathbf{x})-\frac{\mathbf{P}+\bm{\Theta}}{Q}$ for $i=1,\dots,d$, and using $|Q|=q^{t}$. To prove ($\mathfrak{E}$.2), using the ultrametric inequality, for every $j=1,\dots,m$, we have
$\begin{split}\bigg|Qf_{j}(\mathbf{x})-\sum_{i=1}^{d}\partial_{i}f_{j}(\mathbf{x})\big(Qx_{i}-P_{i}-\Theta_{i}\big)-P_{d+j}-\Theta_{d+j}\bigg|&\leq\max\Big\{|Qf_{j}(\mathbf{x})-P_{d+j}-\Theta_{d+j}|,\\ &\max_{i=1,\dots,d}|\partial_{i}f_{j}(\mathbf{x})|\cdot|Qx_{i}-P_{i}-\Theta_{i}|\Big\}.\end{split}$
Since $\big\|\mathbf{f}(\mathbf{x})-\tfrac{\mathbf{P}+\bm{\Theta}}{Q}\big\|<q^{-(s+t)}\,\,\text{and}\,\,|Q|=q^{t}$, it follows that $|Qf_{j}(\mathbf{x})-P_{d+j}-\Theta_{d+j}|<q^{-s}$. Combining this and ($\mathfrak{E}$.1), we get
$\max\left\{|Qf_{j}(\mathbf{x})-P_{d+j}-\Theta_{d+j}|,\max_{i=1,\dots,d}|\partial_{i}f_{j}(\mathbf{x})|\cdot|Qx_{i}-P_{i}-\Theta_{i}|\right\}<Mq^{-s}.$
This proves ($\mathfrak{E}$.2).
@@
Suppose that $\mathbf{x}\in U$ and $(\mathbf{P},Q)\in\mathcal{R}^{n}\times(\mathcal{R}\setminus\{0\})$ satisfy the hypothesis of Lemma 4.1. Let $\mathbf{x}_{0}\in B\left(\mathbf{x},q^{-\frac{s+t}{2}}\right)$, where $\mathbf{x}_{0}=(x_{1,0},\cdots,x_{d,0})\in U$. If $t>s>0$, then
  1. (i)
    $|Qx_{i,0}-P_{i}-\Theta_{i}|<q^{\frac{t-s}{2}}$ for $i=1,\dots,d$; and
  2. (ii)
@@
    $\left|Qf_{j}(\mathbf{x}_{0})-\sum_{i=1}^{d}\partial_{i}f_{j}(\mathbf{x}_{0})(Qx_{i,0}-P_{i}-\Theta_{i})-P_{d+j}-\Theta_{d+j}\right|\leq Mq^{-s}.$
@@
Let $\mathbf{x}_{0}=(x_{1,0},\cdots,x_{d,0})\in U$. By Lemma 4.1, we see that the vector $\mathbf{x}$ satisfies ($\mathfrak{E}$.1) and ($\mathfrak{E}$.2). Hence,
$|Qx_{i,0}-P_{i}-\Theta_{i}|\leq\max\left\{|Qx_{i}-P_{i}-\Theta_{i}|,|Q|\cdot|x_{i}-x_{i,0}|\right\}\leq\max\left\{q^{-s},q^{\frac{t-s}{2}}\right\}=q^{\frac{t-s}{2}}.$
Also, for every $j=1,\dots,m$, it is easy to see that
$\begin{split}\left|Qf_{j}(\mathbf{x}_{0})-\sum_{i=1}^{d}\partial_{i}f_{j}(\mathbf{x}_{0})(Qx_{i,0}-P_{i}-\Theta_{i})-P_{d+j}-\Theta_{d+j}\right|\\ \leq\max\Bigg\{\left|Qf_{j}(\mathbf{x})-\sum_{i=1}^{d}\partial_{i}f_{j}(\mathbf{x})(Qx_{i}-P_{i}-\Theta_{i})-P_{d+j}-\Theta_{d+j}\right|,\\ |Q|\cdot\left|f_{j}(\mathbf{x}_{0})-f_{j}(\mathbf{x})-\sum_{i=1}^{d}\partial_{i}f_{j}(\mathbf{x})(x_{i,0}-x_{i})\right|,\\ \max_{i=1,\dots,d}|\partial_{i}f_{j}(\mathbf{x}_{0})-\partial_{i}f_{j}(\mathbf{x})|\cdot|Qx_{i}-P_{i}-\Theta_{i}|\Bigg\},\\ \leq\max\left\{Mq^{-s},q^{t}M\left(q^{\frac{-t-s}{2}}\right)^{2},Mq^{-s}\right\}=Mq^{-s},\end{split}$
in view of ($\mathfrak{E}$.2), ($\mathfrak{A}$.4), and Taylor’s formula (2.2). ∎
4.2. Analysis on the Generic Part
Recall that $\mathcal{K}_{2(n+1)}$ denotes the extended field $\mathcal{K}\left(T^{\frac{1}{2(n+1)}}\right).$ We first define the generic part. To do so, we first define the diagonal flow $g_{s,t}\in\operatorname{SL}\left(n+1,\mathcal{K}_{2(n+1)}\right)$ as
$g_{s,t}{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\operatorname{diag}\left\{\underbrace{T^{\frac{(d+2)t+2s}{2(n+1)}},\dots,T^{\frac{(d+2)t+2s}{2(n+1)}}}_{m},\overbrace{T^{\frac{(d+2)t+2s}{2(n+1)}-\frac{s+t}{2}},\dots,T^{\frac{(d+2)t+2s}{2(n+1)}-\frac{s+t}{2}}}^{d},T^{\frac{(d+2)t+2s}{2(n+1)}-s-t}\right\}.$
For $\mathbf{x}\in U$, denote by $J(\mathbf{x})$ the Jacobian matrix of $\mathbf{f}$ at $\mathbf{x}$, i.e.,
$J(\mathbf{x}){\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\left(\frac{\partial f_{i}}{\partial x_{j}}(\mathbf{x})\right)_{1\leq i\leq m,1\leq j\leq d}.$